OpenStudy (anonymous):

Solve: −negative 1 over 2 x + 1 = −x + 8

OpenStudy (anonymous):

Is the problem: \[\frac{-1}{2x}+1=-x+8\]? Your wording is confusing

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OK. Multiply through by 2x. \[-1+2x = (-x)(2x)+8(2x)\]\[-1+2x = -x^2+16x\]\[x^2-14x-1 = 0\] And use the quadratic formula